LASCO/EIT Operations
Detailed Status Reports During Recommissioning Activities

   SOHO S/C Recovery Activities

   SOHO News..."What's New"

Feb 18th 99:

After a brief delay in observations,SOHO is back in Medium Rate Telemetry and is observing with EIT 171A images.

At 13:55 UT this morning the operations by the FOT went smoothly and submode 6 telemetry was resumed. The LEB was working properly and EIT was allowed to send a single command to get images for pointing corrections at 14:46 UT. Click here for the first image showing the pointing. We are currently rolled -120 degrees so the North Pole appears to be the South West. Sun center is approx. 1 arc minute off which will be adjusted today.

SOHO will remain in this roll position until March 1st scheduled correction date.

Feb 2nd 99:

SOHO Is In Normal Mode And Back In Medium Rate Telemetry!

All is well and the LASCO/EIT instruments are fine. Calibrations will continue through the next 2 days then regular observations will resume.

Be sure to check the LASCO daily status page for times and events.

Feb 1st 99:

SOHO Is Out Of ESR Mode!

S/C Operations team did a great job over the weekend. Recovery now being performed to reconfigure from a 'Warm-Startup' late Saturday evening.

Scheduled transition to medium rate telemetry is currently set for Tuesday Feb 2nd and NRT may be possible late that day.

Check the SDAC Home Page for a look at a well deserved FOT "Thanks".

Jan 29th 99:

The S/C recovery and transition activities have begun.

At 15:00 UT today the FOT started the transition procedures. Click Here To See The Near-Live Updated Schedule of the ESR Escape Maneuvers being performed. You will see the planned maneuver schedule for the next two days as well.

Jan 26th 99:

Preparations for the instrument commissioning are being planned.Transition to RMW (roll maneuver wheels) mode as early as Saturday Jan 30th and transition to normal mode by Monday Feb 1st.

The LASCO and EIT cameras will be going through full calibration sequences after the initial instrument status has been determined. EIT will be delaying the planned CCD bakeout for the time being.

The time-line for instrument medium rate telemetry is Tuesday February 2nd. Normal observations and cadence will resume after the calibrations are complete.

Station Keeping maneuver #3 while in ESR mode executed today. Correction of delta-V of 3.8 m/s with fuel usage of 4 kg.

S/C still holds in ESR safe mode.

Jan 19th 99:

Station Keeping maneuver #2 while in ESR scheduled for 18:00 UT today.

NASA/ESA engineers have developed a new program for the delta-V correction which will not oppose the ESR thruster firings and will conserve the S/C fuel usage for the maneuver.

S/C still holds in ESR safe mode.

Jan 12th 99:

Thursday Jan 7th was the first station keeping while in ESR mode. The fuel usage for the maneuver was estimated to be 3.5 Kg but consumed 9 Kg total due to the competition from the ESR thruster firings. There is now 170 Kg fuel on-board SOHO.

Flight Operations Team had a very busy weekend.

Tests were run and successfully completed to reduce the ESR thruster firings from 15 per 1 hour time-span to 6 per hour. While the ESR gives numerous thruster pulses automatically, the objective is from the ground to give a minimal opposing thruster pulse to 'slow' the return of the S/C before the ESR would send its large opposing pulse. This test cut fuel usage by one-half and is currently implemented.

To transition back to RMW mode, one option is to lock the guide star tracker. Two tests were done over the weeekend and successfully gained lock and tracked for several minutes during a low roll activity period. Once RMW mode is obtained, reaction wheels and star tracker mode will be used for control which will eliminate the thrusters and fuel consumption. Software changes will be made to eliminate the need for gyros.

S/C still holds in ESR safe mode.

Dec 21st 98:

Mission Interuption
No Observations While S/C Holds In ESR Safe mode

Dec 18th 98:

The Z micrometer voltages reversed at 03:15 UT last night. The DACS are now receiving valid readings and is behaving properly. The increase of the temperature in the COB is a factor but as of yet the reasons for the strange activities in the FP are not fully understood.

Plots of the OCC 1,2,3 , the Z micrometer and the changes of the PZTs below:

DACS Values

  •  Dec 17 00:00 UT - Dec 18 10:00 UT
  •  Dec 18 00:00 UT - 06:00 UT
  • with better resolution.
  •  Z Micron & Digital Offset Dec 18 17:00-17:30 UT
  •  Z Micron & Digital Offset Dec 18 17:10-17:30 UT
  • with high resolution.

    PZTs Voltages

  •  Dec 17 00:00 UT - Dec 18 10:00 UT
  •  Dec 18 00:00 UT - 06:00 UT
  • with better resolution
  •  PZT 3 Dec 18 17:00-17:30 UT
  •  PZT 3 Dec 18 17:12-17:30
  • with high resolution.

    Dec 17th 98:

    The DACS is still reading invalid values as seen in the positive (+) value readings from the Z micrometer voltages. The temperatures of the COB are elevated +7.5 degrees from nominal to help with the electronics and mechanical elements of the FP.

    Plots of the OCC 1,2,3 and Z micrometer below:

  •  Dec 16 00:00 UT - Dec 17 09:00 UT

  • Dec 15th 98:

    Manuall adjustments of the DACS were made. Plots of the OCC 1,2,3 and the Z micrometer below for observation and analysis.

  •  Dec 15 14:00 UT - 22:00 UT
  •  Dec 15 14:00 UT - 22:00 UT
  • with better resolution.

  •  Dec 15 16:00 UT - Dec 16 17:00 UT

  • Dec 14th 98:

    Plots for the Occ 1,2,3 and Z micrometer voltages:

  •  December 14 04:00 UT - 16:00 UT
  • with better resolution.
  •  December 13 00:00 UT through Dec 14 16:00 UT
  •  December 14 00:00 UT through Dec 15 14:00 UT

  • Dec 13th 98:

    FP still in servo mode. Adjustments will made to put the DACS range into the 'zero null-point' with the PZTs early this week.

    Plots for the OCC 1,2,3 and Z micrometer voltages below.

  •   December 12 00:00 UT through Dec 13 21:00 UT

  • Dec 11th 98:

    A LEB reboot at 03:50 UT due to the EIT sector wheel hang during the 01:00 UT synoptic set put the FP back in Servo mode and the PZT voltages at full (112 volts) and will remain through the weekend. The COB temperatures reset to the + 5.5 degrees above nominal for the FP analysis and FeXIV/Clear images at a 2 hour cadence.

    FP plots below show the OCC 1,2,3 and Z micrometer voltages for this period:

  •  December 10, 1998 00:00 UT through December 11 17:30 UT
  •  December 10 00:00 UT through December 11 17:30 UT
  • with better resolution.

    Dec 9th 98:

    Monitoring of the FP while in reduced mode at the 'zero null-point' continues.

    Plots of the OCC 1,2,3 and the Z micrometer voltages below.

  •  December 9, 1998 00:00 UT-22:30 UT
  •  December 7, 1998 00:00 UT through December 9, 1998 17:00 UT

  • Dec 7th 98:

    Late Friday evening on Dec 4th, the DACS were manually adjusted to as close to a 'zero null-point' as possible and left over the weekend for observations with images at the 2 hour cadence.

    We saw many changes in the Z micrometer and OCC (Optical Control Channel), 1, 2, and 3 voltages. Plots are made for the time periods of interest:

  •  December 4, 1998 00:00 UT-04:00 UT
  •  December 5, 1998 00:00 UT-December 7, 1998 11:50 UT
  •  December 5, 1997 00:00 UT-December7, 1998 11:50 UT
  • with reduced min/max values.
  •  December 7, 1998 00:00 UT-19:50 UT
  •  December 7, 1998 15:00-20:30 UT
  • with better resolution.
  •  December 7, 1998 20:00 UT-December 8, 1998 12:00 UT
  •  December 8, 1998 15:00 UT-21:00 UT
  • with better resolution.
  •  December 5, 1998 00:00 UT through December 8, 1998 20:00 UT

  • Dec 4th 98:

    The FP has shown change. The attempts of determining the cavity length and DACS range by way of setting DACS to zero while PZT voltage remained at full, changed the Z Micrometer from the previous reading of +8.2 volts to -9.0 volts. See Plot

    This is a saturation of the Z micometer in the 'extreme' opposite 'long-side' of the cavity. The cavity length and the range values will be determined along with the 'null-point' in the Z Offset to set the Range DAC as close to zero as possible. This will give a 'readable' range factor.

    Continued analysis and FeXIV/Clear images are being taken at a 2 hour cadence while still the LASCO COB heater zones being increased by +5.5 degrees.

    Dec 2nd 98:

    SOHO Review Board Hearings begin today.

    FP servo 'aliveness' testing performed. An attempt to calculate the cavity length using the FP electronics by adjusting the set points of the DACS to the 'extreme' ends of zeros to highs.

    The three PZTs are operating at full 112 volts while the DACS were dropped to zero. In hopes to align a range between the mirror plates in high and DACS in low, we will see with the known change in electronics and the error to the full cavity length of 43mm, + or - 6mm.

    The LASCO heater zones were increased + 5.5 degrees to again 'flex' the LASCO COB and restart the 'walk-off' correction. FeXIV/Clear images at a 63 second exposure with a 2 hour cadence continue for analysis.

    Nov 24th 98:

    Return to nominal mission is scheduled for December 2, 1998.

    C1 FP testing showed a ceiling to the mirror plate correction recovery. We will attempt to manually adjust the plates to parallel. A manual adjustment will be determined by aligning the plates and taking the images. PZT #3 was adjusted to zero volts while PZTs 2 and 3 remain at 100 volts. We will take images in FeXIV/Clear at 63 second exposures on a 2 hour cadence for analysis.

    Nov 20th 98:

    Instrument power-on at 20:30 UT. Precautions were taken to safe for the Leonids Meteor Shower interuption. The instruments were powered-off November 15th to protect on-board electronics.

    C1 is continuing the FeXIV/Clear door closed program with a 1 hour cadence. We have found using the 3 hour cadence for the days of Nov 9th-12th, the PZTs are "on their own" recovering. The images are showing a linear correction to the mirror plate 'walk-off' with a +25% improvement from the images of reference dating Oct 22nd. We are increasing the cadence and will study further the changes that are taking place in the FP.

    Nov 9th 98:

    A C1 FP program is started taking door closed FeXIV / Clear images at a 63 second exposure. These images are taken every 3 hours and will continue to 19:00 UT Thursday November 12th.

    The images are going to be used in animation sequence to determine if the mirror plates are seeing change in succession as the PZTs are kept at constant voltages while in servo mode.

    Nov 5th and 6th 98:

    The C1 Fp testing continued. PZT #3 movement range testing took reference images while in servo mode and while making 33% voltage decreases on PZT #3 only, we developed the distance of travel and determined the movement range. We will use this calculation along with the current wedge defect walk-off and find the proper mirror plate displacement.

    Oct 29th 98:

    C1 Fp testing continued with the increasing of the thermal cycle and a 5 DegC temperature increase from the nominal operating environment in the COB (Coronagraph Optical Box). The duration of the temperature increase was 24 hours respectfully, Oct 29, 22:32 UT to Oct 30, 22:52 UT. Minor changes can be seen in C2 and C3 data due to this temperature increase.

         COB Zone

      Zone Set Point

      Zone Temp

      Zone Set Point

      Zone Temp

         LASCO 1             110            15.5             97             19.7 
         LASCO 2             106*                    16.8            96            20.0
         LASCO 3             109*            15.8            95            20.5
         LASCO 4             110            15.5            94            20.8
         LASCO 5             110            15.5            94            20.8
                    *The  Nominal Set Point Values are 110 for all zones.
                       Zone 2 and 3 appear reduced due to C1 door being closed.

     The thermal cycling was an attempt to correct the current wedge defect of the misalligned mirror plates by intentionally causing flex in the box, then making the necessary adjustments in the PZTs before the cool-down begins. Reference images were taken before and after in FeXIV/Clear filters to verify settings. 

    Continued evaluation and preparations being made for the next test to begin Wednesday November 4, 1998.

    Oct 23rd 98:

    Weekend observation plan assembled to complete the C2 and C3 calibrations with the doors open. Four step polarizer sequences using the clear, +60 degree, zero degree, and -60 degree polarizers with various color filters for both cameras. C2 polarizers in BLUE, DEEP RED, and ORANGE filters, and a normal image with H-ALPHA/CLEAR. C3 polarizers in CLEAR, DEEP RED, ORANGE , and BLUE, and normal image in IR/CLEAR.

    The C2 polarized brightness images are perfect and will be of needed use in the upcoming 40 hour SPARTAN-201-5 mission for calibration scheduled for November 1st.

    C3 is found to have pressure on the zero degree polarizing filter. This is dissapointing, but having the +60 degree and -60 degree filters correct, we can get the appropriate values for the defective zero degree filter and still get accurate sums in the images.

     C1 FP testing and commissioning to resume Thursday Oct 29th.

    Oct 22nd 98:

    Continued testing of the C1 FP was a defined set of PZT movements from the program run on October 21st. The PZTs again responded to the voltage changes as expected, but the low to high voltage changes were alternated for each PZT which singled the high voltage position.

    These images are quite similar to the previous evenings images.

     Oct 22 19:33:15 UT

    Oct 22 19:57:06 UT

    Oct 22 20:16:01 UT

    Oct 22 20:29:12 UT

    Oct 22 20:44:53 UT Notice the extension of the walk-off. This is showing the wedge defect to a different plane than noticed earlier.

     This M-Peg of Oct 22nd shows the 5 images together. This helps to notice the change.

     We have seen the movement now from all three PZTs. We have also determined the cavity length is outside of the electronic range of the cap mics, which also the wedge defect is outside the range of the PZTs to compensate for it.

    The next step will be to determine the actual cavity length. 

    Oct 21st 98:

    Todays testing of the C1 FP was designed to check for PZT movement and provide wedge defect direction and magnitude calibration data. PZT voltages were permutated so that one PZT was at +100V and 2 PZT's were at 0V. The linear extension of the PZT's as a function of applied voltage can be calculated from the image data. The magnitude and direction of the wedge defects can be calculated from the spacing of the arcs in the 'walk-off' images. The 'arcs' are multiple reflections of the inner edge of the occulter hole. Changes in the polar angle and magnitude of wedge defect from image to image indicate that at least some of the PZT's are responding to applied voltage changes.

    As you notice in the image of Oct 21 23:14:06 UT the wedge polar coordinates angle is about 120 deg.

    The image of Oct 21 23:23:11 UT shows the wedge polar coordinates angle is slightly increased after a voltage step increase.

     The image of Oct 21 23:54:44 UT shows a drastic polar coordinate change to 60 degrees after voltage increase.

     The image of Oct 22 00:17:30 UT takes the walk-off back to the polar coordinate angle of 120 degrees.

     Then after going to full +112V the image of Oct 22 00:31:07 UT saw the return of the image from Oct 20th. This was an expected movement of the PZT's and from here we happily called it a night.

    Oct 20th 98:

    Continued testing of the C1 FP revealed response from the PZT's. The C1 camera mechanisms were tested and found to be working properly. The first door closed normal image from C1 since reacquisition shows blur due to the plates slightly misalligned. The test results are being analyzed and continued programs will be run to determine the status of the FP tomorrow.
    C1, C2, C3 Dark images were taken at 0, 60, and 300 second exposures. These calibrations will be compared to the previous images to show any thermal changes from the power-on period.

    Oct 19th 98:

    Finished the schedule from the two previous days. EIT full set of calibations were performed. This is the same set as is done after bakeout. It also verified that the light leaks in the prefilters have not gotten worse.

    The checkout of the FP started by sending commands to set the PZT voltages to specific values while in reduced capability mode. In this mode, no change in the capacitance microcrometers were detected, while the error signal increased. Checkout will continue.

    An analysis of the C3 door calibration images reveals a 1-2% change in the C3 polarizer transmittances in the small region where the calibration window appears. The inability to do the C3 polarization calculation can not be explained as a simple change in the polarizer transmittance. An analysis of the calibration lamp images shows a considerable change in the 0 degree polarizer from before the extreme cold soak.

    Oct 18th 98:

    Continued the schedule from yesterday.

    Oct 17th 98:

    EIT normal imaging. C1/C2/C3 doors are closed. C2 and C3 calibration lamp images through all filter and polarizer combinations. C2 and C3 images of the door through all filter and polarizer combinations.

    Oct 16th 98:

    No checkout due to the S/C momentum management maneuver. EIT took images through the maneuver. The current roll position is not zero, but is about 4 degrees away from solar north. A schedule for the weekend to take C2 and C3 door closed calibration images will be put on-board along with EIT normal images. The EIT calibration sequence is currently planned for Monday.

    Oct 15th 98:

    The checkout of LASCO/EIT continues to be very successful. Absolutely no problems were encountered during the checkout of the C2 telescope system. Door and shutter calibration lamp images were taken with the door closed. Then the C2 door was opened and the corona was imaged first in H-alpha, and then with the normal orange filter. The images are excellent. No image degradation is perceived. The pointing appears to be shifted a little, but this might be due to thermal bending of the box, since the door was just opened and the C1 door remains closed.

     After the initial images, polarization sequences in C3 and C2 were taken, followed by a synoptic sequence in EIT. Evaluation of the pB images shows no problems with the C2 images, but does show a problem when we calculate pB for C3. We will be performing a detailed calibration of the filter transmittances this weekend after which we may understand the problem in calculating the C3 pB.

    Evaluation of the noise and bias levels in EIT and C3 images, shows no change in those parameters.

    The C2 and C3 doors were closed in preparation for the spacecraft maneuvers tomorrow.

     The continued operation in EIT, C3, and C2 is giving us more confidence in the stability of the instrument. In summary, the LEB, EIT, C3 and C2 are all functional. The only mechanisms that have not been checked on the three telescopes are the EIT door, the C2 focus and the C2 IOCS. Absolutely no errors have been detected in the operation. Detailed calibration activities will be proceeding during the next days to weeks. The functional checkout of the C1/FP system will hopefully take place next week. 

    Oct 14th 98:

    The checkout of LASCO/EIT continued today beginning at about 1pm EST (1700UT). The NRT was much shorter than yesterday. We started by taking a full synoptic sequence in the four EIT sectors. Then after a pause for FOT work, we continued the checkout. After completing the EIT filter wheel check, we concentrated on checkout of the C3 telescope with an aim to opening the C3 door if the checkout went as smoothly as yesterday. We performed a full functional checkout of C3, found that the door calibration window image was producing counts that were about what we had seen before and decided to open the door. The door opening went very smoothly. We then took a 4x4 on-chip binned image in H-alpha to make sure that the optical alignment was still intact. Then we took a normal clear/clear image of the corona. MARVELOUS. We then decided to put a schedule of EIT and C3 running through the night, after watching the progress of the schedule for about 90 minutes. 

    Oct 13th 98:

    The checkout of LASCO/EIT began today at about 1pm EST (1700UT). Much more was achieved today than expected because the checkout went so smoothly.

    Checkouts of the LEB and EIT have been completed and they work perfectly! All five processors in the LEB and EIT along with the four in the COB are working perfectly, with no hiccups at all in the inter-processor (or intra) communications, etc. ICE and OBE booted up out of EEPROM, showing that it survived and giving us the revised versions of those two modules (the others were the same as in bipolar ROM). The thermal control of the COB and EIT has been switched to internal (LEB) control and is holding the temperature zones to the proper values. With this new thermal environment, the internal temperatures will probably shift over the next few hours.

    EIT images in all four sectors have been taken and are absolutely beautiful, showing the benefit of having been "baked out" at a higher temperature than we can achieve normally. This is the result of the amazing achievement by the SOHO recovery team. THANK YOU. EIT camera functions (ie, read out port, on-chip summing, sub field readout, heaters) were verified. The sector wheel worked normally. The checkout of the EIT filter wheel was deferred until tomorrow.

    Dark images (0, 60, 300 seconds) from all telescopes have been acquired. As expected, some shifts in the bias values are noted, but more analysis will need to be made. Qualitatively, all the CCDs appear to be in good shape, with no extra hot or dark pixels detected. The temperatures of the CCDs were about 3 degrees warmer than they had been, but still very cold. This might be due to the camera temperatures being slightly warmer than they will be when they get to the normal operating condition.

    All shutters and calibration lamps in the EIT and COB are working properly. The C2 filter and polarizer wheels are also OK. The FP status is consistent with the concept that the piezoelectric crystals are not working, but we have not done the testing to prove it.

    The pointing of the COB as determined by the boresighters is about where it was in June. The BS-Y boresighter continues to show a slight drift. EIT shows that the pointing is also where it was. The roll of about 50 degrees will be corrected at the next station keeping manuever, currently this friday.

    Most of the testing was done today in telemetry submode 1 (5.2 kbps) but then it was switched to submode 5 (7.9kbps). No problems at all in either submode. Submode 6 will be uploaded soon, giving us nearly 15.8 kbps.

    Testing will continue on the COB tomorrow after a full set of EIT images is taken.